OpenStudy (anonymous):

What is the solution of the system? 3x – 4y = –24 x + y = –1

OpenStudy (pfenn1):

Solve the second equation for x in terms of y, plug it into the first equation for x and solve for y.

OpenStudy (anonymous):

x = -1 -y 3(-1-y) - 4y = -24 -3 - 3y - 4y = -24 -7y = -21 y = 3 therfore x = -4